# =COLI-manual3.qry Cross-lingual Cross-temporal Summarization: Dataset, Models, Evaluation

### v1

@@ 3.4 Historical Language Divergence
Languages evolve due to social, cultural and linguistic pressure and this poses challenges to the summarization of historical texts.
@@
The distribution of publication year is shown in Figure 2. HISTSUMM documents (in red) distribute evenly between the time period of 1650–1800. CLCTS hEn-De (in gray) contains more documents from the late 1800s and CLCTS hDe-En (in green) from the early and mid 1800s. The documents from CLCTS hEn-De which date back to the 1600s are plays from Shakespeare. For both CLCTS datasets, the majority of historical documents come from the 1800-1900s and comparatively HISTSUMM has overall older documents than our CLCTS corpus.
@@
Divergence in syntax Length of Sentences The analysis in Table 2 (column 5) reveals that historical German documents in HISTSUMM and CLCTS corpus contain longer sentences compared to other datasets, where HISTSUMM hDe-De has a mean sentence length of 36.3 tokens and CLCTS hDe-En/hDe-De dataset has 25.4 tokens per sentence while modern German documents from Wikilingua and MLSUM have maximum 20 tokens per sentence on average. This coincides with the finding of decreasing sentence length over time (Rudnicka, 2018), considering that HISTSUMM documents are the oldest. Slightly surprisingly, this is not true for the English documents, where the mean sentence length of historical English documents (CLCTS hEn-De/hEn-En) is almost the same as Wikilingua and is even shorter than documents from news genre CNN/Daily Mail. This hints at a genre effect, besides the temporal effect: fiction, especially plays, may have shorter sentences than news and other non-fiction genres due to the higher occurrence of dialogues and two-word sentences (Rudnicka, 2018). In our CLCTS corpus, unlike CLCTS hDe-En/hDe-De containing only fairytales, historical English documents contain plays. When we exclude plays from Shakespeare, the mean sentence length of the CLCTS hEn-De document increases by 4 tokens per sentence to 20.3, which is closer to the value of CNN/Daily Mail. The factual summaries from CLCTS hEn-De/hEn-En in column 6 also contain longer sentences than their fictional documents for similar reasons.
Dependency Distances To further quantify syntactic changes, we analyze the dependency parsing trees for sentences from documents in all datasets and compute the mean dependency distance (MDD) over the entire corpus Liu (2008); Liu, Xu, and Liang (2017).
This measure can reflect the syntactic complexity of sentences. We use stanza (Qi et al., 2020) for parsing the sentences. We plot the
MDD
@@
in Figure 3 where we observe a downward trend overall. The first four documents from the left are in German and the last three documents are in English (we use dataset plus language e.g., CLCTS hDe to represent the documents here). The downward trending line reveals an overall higher MDD of German sentences than that of English sentences. If we take a closer look by language, we observe: (i) For German (the first four points from the left), historical sentences (in red star) have a higher value for MDD than that of the modern German text, which suggests more syntactic complexity of historical documents. We especially notice a higher MDD from CLCTS hDe documents (fiction) than that from HISTSUMM hDe (news), which coincides with the finding of longer dependencies in the imaginative genre than those in the informative genre from Wang and Liu (2017). The modern sentences from MLSUM are also more complex syntactically compared to Wikilingua. (ii) For English (top-3 right items), this is not entirely true. News documents from CNN/Daily Mail have again a longer MDD than that of Wikilingua. However, they also have a longer MDD than CLCTS hEn.
This may be because of the combined effect of (a) shorter sentences from CLCTS hEn (also from Wikilingua En) as discussed above, (b) more conversational texts with shorter dependencies from CLCTS hEn documents (plays and short stories) Wang and Liu (2017) and (c) CLCTS hEn documents mostly from the late 1800s which share more similarities to modern English.
@@
Divergence in lexicon The vocabulary used in historical and modern language varies greatly due to spelling, morphological and semantic changes. For example, in our historical German dataset, the letter “ß” often represents “ss” in modern texts (e.g., “daß” vs. “dass”). Omitting “h”, “t” or “e” is another type of spelling change such as “Thüre” (door, now “Türe” or “Tür”), “thun” (to do, now tun), “Todt” (death, now “Tod”) and “gieng” (passive tense of to go, now “ging”).
Such changes in German are a result of two spelling reformations, one in 1901 and one in 1996.
In English, we can also observe changes in lexicon such as the word “you”
written as “thou” and the morphological change of the second person “will” as “wilt”. To quantify these divergences, we compute the Jaccard similarity (Niwattanakul et al., 2013) of lexicons between documents and their summaries in Table 4. Historical German has lowest lexical overlap with modern German, with a Jaccard score of 0.1 for HISTSUMM and 0.19 for CLCTS hDe-De.
The pattern is similar for English.
@@
Semantic shift is another source of language change (Eger and Mehler, 2016; Hamilton, Leskovec, and Jurafsky, 2016). For example, the word “mistress” in Shakespeare’s work has a similar meaning as “sweetheart”. The “car” mentioned in Shakespeare’s work (e.g., “Phoebus’ car”) refers to a special kind of chariot. Another example is the adjective “gay” which is used in the sense of “glad and cheerful”, e.g., “The great city in which he dwelt was very gay, for every day strangers visited the town” (The Emperor’s New Clothes by Christian Andersen in 1888) . Similarly, the meaning of the word “knave” degenerates from “servant” in the story of brother Grimm to the modern meaning “a dishonest man” over time.
@@ 4 Experimental setup
Our experiments consist of three methods: pipeline methods based on extractive summarizers, e2e transformer-based methods with intermediate task finetuning and zero-shot ChatGPT with different prompt strategies. For finetuned methods (extractive and abstractive), we use 5-fold cross-validation to obtain reliable results for all CLCTS datasets. For ChatGPT, we query summaries for all CLCTS datsets as well to make sure the results are comparable.
4.1 Pipeline based on extractive summarizer
@@
For extractive summarization, we first use an extract-then-translate pipeline. We choose MemSum (Gu, Ash, and Hahnloser, 2022) as our model, which utilizes multi-step episodic Markov decision processes to extract important sentences from a given text. We use 5-fold cross-validation to train and test MemSum with our monolingual dataset. We limit the maximal extracted sentences per text to 25 according to the average length of summaries in Section 3. During inference, we sort the sentences according to the original text’s order, then translate the extracted sentences to the target language using a multilingual machine translation model developed by Facebook (M2M) (Fan et al., 2021). We additionally experiment with historical text normalization before translation. We use the normalization method Norma666https://github.com/comphist/norma (Bollmann, 2012). We choose the combined version which leverages Lookup, Rule-based and Distance-based normalization which is claimed to perform well for both English and German (Bollmann, 2019).
To tackle input limitations of ChatGPT (e.g., for GPT-3.5-turbo a total of 4096 tokens are shared between prompt and completion777https://help.openai.com/en/articles/4936856-what-are-tokens-and-how-to-count-them), we build the third retrieve-then-summarize pipeline, where we use MemSum as a key sentence retriever and input the retrieved sentences to ChatGPT and request summarization in the target language. For this experiment, we limit the maximal extracted sentences per document to 100 according to the input limits of ChatGPT.
4.2 Transformer-based e2e models
We modify pretrained models together with intermediate task
@@ 4.2 Transformer-based e2e models
finetuning to build e2e models. We propose a modified multilingual mBART50 model pretrained for machine translation Tang et al. (2020) with an efficient attention mechanism taken from LED (Beltagy, Peters, and Cohan, 2020) as model architecture. This architecture is suitable for multilingual long sequences as it benefits from the cross-linguality of mBART and the efficiency of processing long documents from LED. We refer to this model as mLED.
@@
We first provide baselines of mLED finetuned only within the CLCTS corpus: i)
for baseline 1, we finetune the model only with the target task; e.g., for the model in direction hDe-En, we only finetune mLED with our training set of the same direction. We refer to this model as Base.
ii)
For baseline 2 (baseline 1 + historical machine translation), we utilize our historical translation dataset. We first finetune with the translation task and then with the target task, which we call Historical MT.
iii) Finally, for baseline 3, we finetune with datasets from both directions. We add prefixes to indicate the directions. We refer to this model as .
Intermediate task finetuning
Apart from the final finetuning step with our target CLCTS task, we conduct intermediate task finetuning using external sources including MLS, CLS, and CTS. Table 5 depicts the detailed design. We use 5-fold cross-validation to obtain reliable results. Due to resource limitations, our mLED models have a maximum of 4,096 input tokens and the number of tokens for the generated summary ranges from a minimum length of 100 tokens to a maximum length of 512 tokens.
@@
(b) Direction hEn-De |
4.3 CLCTS with ChatGPT
@@
via the OpenAI API to query the summary. We set the temperature to 0.7 to allow for some randomness. We adopt (i) a pipeline translate-then-summarize prompt as well as a (ii) e2e prompt. We set prompt lines to the same language as the target output since ChatGPT works better in this setting according to our initial experiments.888During our first round of data collection all the prompt lines were in English. We noticed that ChatGPT tends to output summaries in the wrong target language, especially when the summary is expected to be in German.
We truncate the texts that exceed ChatGPT input limits.999According to our experiment with the GPT2 tokenizer, we find that on average our German texts have 1.73 tokens per word (which is higher than the suggested 1.3 tokens per word ratio from OpenAI) and English 1.07 tokens per word. Therefore, we limit the German text to 2048 words and English to 3000 words. Since the 4096 tokens limits from GPT-3.5-turbo are shared between both prompts and completions (outputs), we need to leave some tokens for the output summary.
@@
Experiment | Src-Trg | Prompt |
e2e | hDe-En |
Please summarize the following text in English : [text] |
e2e | hEn-De |
Bitte fasse den folgenden Text auf Deutsch zusammen: [Text] |
e2e (Title) | De-En |
Please give me the summary of the story [Title] written by [Author]. |
e2e (Title) | En-De |
Bitte gebe mir die Zusammenfassung der Geschichte [Titel] von [Autor]. |
pipeline | hDe-En |
Please first translate the following text into English and summarize the translated text: [Text] |
pipeline | hEn-De |
Bitte übersetze zuerst den folgenden Text auf Deutsch und fasse den übersetzten Text zusammen: [Text]. |
@@ 4.4 Evaluation
To identify the best model, we utilize multiple recently popular reference-based metrics. We also conduct a multi-phase
human evaluation to evaluate the model and test the effectiveness of the resulting metrics. During the multi-phase evaluation, we give feedback to annotators after each phase for them to better align with the instructions.
Automatic Evalution
We take the ROUGE-variant Lin (2004) which measures the lexical overlap by matching the texts through computing the unigram overlap (ROUGE-1), and ROUGE-L, which leverages the longest common subsequence. BERTScore Zhang et al. (2020b) and MoverScore Zhao et al. (2019) focus on semantic overlap, where BERTScore calculates the token similarity and Moverscore utlizes Word Mover’s distance. We also consider generation-based BARTScore Yuan, Neubig, and Liu (2021), NLI-based MENLI Chen and Eger (2023) and discourse-based DiscoScore Zhao, Strube, and Eger (2023). The configurations used for each metric are presented in Table 14 (appendix).
Human Evaluation
We use six ‘expert’ annotators to evaluate all models.
Three of the annotators are female and three are male.
All annotators have high competence in English as they are affiliated with the university: they are Master students (3), postgraduates (1), PhD students (1) and NLP faculty members (1).
Three of them are native German speakers. Thus, the generated English summaries are evaluated by all six annotators and the generated German summaries are annotated by three native speakers.
We utilize the evaluation strategy by Fabbri et al. (2021) where the annotators score the summary from four perspectives: coherence, consistency, fluency, and relevance. The outputs are rated with a score ranging from 1 (worst) to 5 (best) for each dimension and we allow 0.5 increments. After phase 1, we give feedback to the annotators by providing a comparison of their scores with the average scores. The scores that differ substantially are
highlighted and we encourage them to review the differences.
@@
Nonetheless, we make sure that each model is evaluated for at least 4 identical source documents.
ChatGPT evaluation
Since human annotation is costly, we follow related work Gao et al. (2023); Chiang and yi Lee (2023) and test ChatGPT as an evaluator for our CLCTS task. We differ in the following aspects: 1) our CLCTS task is harder to evaluate due to the long length of both reference and generated summary. 2) Our CLCTS task involves multilingual evaluation for both English and German. We give the same instructions to ChatGPT as to the human annotators.
For reproducibility purposes, we set the temperature to 0 to reduce the randomness.
We collect annotations of generated summaries from all models from the documents annotated by humans, resulting in 501 ChatGPT annotations for hDe-En and 384 ChatGPT annotations for hEn-De.

### v3

@@ 3.4 Historical Language Divergence
Languages evolve due to social, cultural, and linguistic pressure and this poses challenges to the summarization of historical texts. In this section, we firstly provide an overview of the publication time of the historical documents from HISTSUMM and CLCTS and secondly, analyze the divergence of historical language from syntax (including the length of sentences and dependency distances) and lexicon perspectives.
@@
The distribution of publication year is shown in Figure 2. HISTSUMM documents (in red) distribute evenly between the period of 1650–1800. CLCTS hEn-De (in gray) contains more documents from the late 1800s and CLCTS hDe-En (in green) from the early and mid 1800s. The documents from CLCTS hEn-De which date back to the 1600s are plays from Shakespeare. For both CLCTS datasets, the majority of historical documents come from the 1800-1900s and comparatively HISTSUMM has overall older documents than our CLCTS corpus.
@@
Divergence in syntax Length of Sentences The analysis in Table 2 (column 5) reveals that historical German documents in HISTSUMM and CLCTS corpus contain longer sentences compared to other datasets, where HISTSUMM hDe-De has a mean sentence length of 36.3 tokens and CLCTS hDe-En/hDe-De dataset has 25.4 tokens per sentence while modern German documents from Wikilingua and MLSUM have maximum 20 tokens per sentence on average. This coincides with the finding of decreasing sentence length over time (Rudnicka, 2018), considering that HISTSUMM documents are the oldest. Slightly surprisingly, this is not true for the English documents, where the mean sentence length of historical English documents (CLCTS hEn-De/hEn-En) is almost the same as Wikilingua and is even shorter than documents from the news genre CNN/ Daily Mail. This hints at a genre effect, besides the temporal effect: fiction, especially plays, may have shorter sentences than news and other non-fiction genres due to the higher occurrence of dialogues and two-word sentences (Rudnicka, 2018). In our CLCTS corpus, unlike CLCTS hDe-En/hDe-De containing only fairytales, historical English documents contain plays. When we exclude plays from Shakespeare, the mean sentence length of the CLCTS hEn-De document increases by 4 tokens per sentence to 20.3, which is closer to the value of CNN/Daily Mail. The factual summaries from CLCTS hEn-De/hEn-En in column 6 also contain longer sentences than their fictional documents for similar reasons.
@@
Dependency Distances To further quantify syntactic changes, we analyze the dependency parsing trees for sentences from documents in all datasets and compute the mean dependency distance (MDD) over the entire corpus Liu (2008); Liu, Xu, and Liang (2017). This measure can reflect the syntactic complexity of sentences. We use stanza (Qi et al., 2020) for parsing the sentences. We plot the MDD in Figure 3 where we observe a downward trend overall. The first four documents from the left are in German and the last three documents are in English (we use dataset plus language, e.g., CLCTS hDe to represent the documents here). The downward trending line reveals an overall higher MDD of German sentences than that of English sentences. If we take a closer look by language, we observe: (1) For German (the first four points from the left), historical sentences (in red star) have a higher value for MDD than that of the modern German text, which suggests more syntactic complexity of historical documents. We especially notice a higher MDD from CLCTS hDe documents (fiction) than that from HISTSUMM hDe (news), which coincides with the finding of longer dependencies in the imaginative genre than those in the informative genre from Wang and Liu (2017). The modern sentences from MLSUM are also more complex syntactically compared to Wikilingua. (2) For English (top-3 right items), this is not entirely true. News documents from CNN/Daily Mail have again a longer MDD than that of Wikilingua. However, they also have a longer MDD than CLCTS hEn. This may be because of the combined effect of (1) shorter sentences from CLCTS hEn (also from Wikilingua En) as discussed above, (2) more conversational texts with shorter dependencies from CLCTS hEn documents (plays and short stories) Wang and Liu (2017), and (3) CLCTS hEn documents mostly from the late 1800s which share more similarities to modern English.
@@
Divergence in lexicon The vocabulary used in historical and modern language varies greatly due to spelling, morphological and semantic changes. For example, in our historical German dataset, the letter “ß” often represents “ss” in modern texts (e.g., “daß” vs. “dass”). Omitting “h”, “t” or “e” is another type of spelling change such as “Thüre” (door, now “Türe” or “Tür”), “thun” (to do, now tun), “Todt” (death, now “Tod”) and “gieng” (passive tense of to go, now “ging”). Such changes in German are a result of two spelling reformations, one in 1901 and one in 1996. In English, we can also observe changes in lexicon such as the word “you” written as “thou” and the morphological change of the second person “will” as “wilt”. To quantify these divergences, we compute the Jaccard similarity (Niwattanakul et al., 2013) of the lexicons between documents and their summaries in Table 4. Historical German has the lowest lexical overlap with modern German, with a Jaccard score of 0.1 for HISTSUMM and 0.19 for CLCTS hDe-De. The pattern is similar for English.
@@
Semantic shift is another source of language change (Eger and Mehler, 2016; Hamilton, Leskovec, and Jurafsky, 2016). For example, the word “mistress” in Shakespeare’s work has a similar meaning as “sweetheart”. The “car” mentioned in Shakespeare’s work (e.g., “Phoebus’ car”) refers to a special kind of chariot. Another example is the adjective “gay” which is used in the sense of “glad and cheerful”, for example, “The great city in which he dwelt was very gay, for every day strangers visited the town” (The Emperor’s New Clothes by Christian Andersen in 1888). Similarly, the meaning of the word “knave” degenerates from “servant” in the story of Brother Grimm to the modern meaning “a dishonest man” over time.
To conclude: (1) Historical German documents from HISTSUMM and CLCTS contain longer sentences with decreasing sentence length over time (temporal effect) while the sentence length of our historical English documents shows barely any difference to modern texts because of the genre effect. (2) German sentences show an overall higher MDD than that of English sentences in our entire corpus due to the joint effect of genre and publication time. (3) The vocabularies of historical and modern languages differ greatly in terms of spelling, morphological, and semantic changes.
@@ 4 Experimental setup
In this section, we introduce the experimental setup of CLCTS. Figure 4 illustrates the model designs (detailed in Section 4.1) and evaluation strategies (detailed in Section 4.2).
4.1 Models
@@
Our experiments consist of three methods: (1) pipeline methods based on extractive summarizers, (2) e2e transformer-based methods with intermediate task finetuning, and (3) zero-shot ChatGPT with different prompt strategies. For finetuned methods (extractive and abstractive), we use 5-fold cross-validation to obtain reliable results for all CLCTS datasets. For ChatGPT, we query summaries for all CLCTS datasets as well to make sure the results are comparable.888During 5-fold cross-validation, we accumulate the test sets from each fold and in this way, we obtain a final test set consisting of all samples. For example, for direction hDe-En with 328 samples, we accumulate output summaries from each test set consisting of 65-66 samples and overall we collect 328 samples after 5-fold cross-validation. In contrast, for zero-shot ChatGPT, we query directly from ChatGPT and collect 328 summaries at once. We summarize model details in Table 5.
Model name | Methods | Type | CV |
MemSum | |||
- | extract-then-translate pipeline with 25 maximal extracted sentences per source document | extractive pipeline | ✓ |
- | extract-normalize-translate pipeline with 25 maximal extracted sentences per source document | ||
mLED | |||
-Base | finetune with CLCTS target task alone | abstractive e2e | ✓ |
-Historical MT | finetune with CLCTS historical translation dataset & CLCTS target task | ||
finetune with CLCTS non-target task & CLCTS target task | |||
mLED | |||
finetune with intermediate tasks listed in Table 6 & CLCTS target task | abstractive e2e | ✓ | |
-MLS-CLS | |||
-MLS-CLS-CTS | |||
ChatGPT as summarizer | |||
-e2e (Title) prompt | prompt in e2e style using only Title and Author | zero-shot abstractive e2e | |
-e2e prompt | prompt in e2e style using source documents | ||
-pipeline prompt | prompt in pipeline style using source documents | ||
- | retrieve max 100 key sentences & prompt in e2e style using retrieved texts | zero-shot abstractive e2e (two-step) |
4.1.1 Pipeline based on extractive summarizer
@@
For extractive summarization, we first use an extract-then-translate pipeline. We choose MemSum (Gu, Ash, and Hahnloser, 2022) as our model, which utilizes multi-step episodic Markov decision processes to extract important sentences from a given text. We use 5-fold cross-validation to train and test MemSum with our monolingual dataset. We limit the maximal extracted sentences per text to 25 according to the average length of summaries in Section 3. During inference, we sort the sentences according to the original text’s order and then translate the extracted sentences to the target language using a multilingual machine translation model developed by Facebook (M2M) Fan et al. (2021). We refer to this model as MemSum-translation. We additionally experiment with historical text normalization before translation. We use the normalization method Norma999https://github.com/comphist/norma Bollmann (2012). We choose the combined version that leverages Lookup, Rule-based, and Distance-based normalization which is claimed to perform well for both English and German Bollmann (2019).101010Norma is recommended by the author when only little training data is available We refer to this model as MemSum-Norma-translation.
4.1.2 Transformer-based e2e models
@@
We modify pretrained models together with intermediate task finetuning to build e2e models. We propose a modified multilingual mBART50 model pretrained for machine translation Tang et al. (2020) with an efficient attention mechanism taken from LED (Beltagy, Peters, and Cohan, 2020) as model architecture. This architecture is suitable for multilingual long sequences as it benefits from the cross-linguality of mBART and the efficiency of processing long documents from LED. We refer to this model as mLED. We use 5-fold cross-validation to obtain reliable results. Due to resource limitations, our mLED models have a maximum of 4,096 input tokens and the number of tokens for the generated summary ranges from a minimum length of 100 tokens to a maximum length of 512 tokens.
@@
We first provide baselines of mLED finetuned only within the CLCTS corpus:
for baseline 1, we finetune the model only with the target task; for example, for the model in direction hDe-En, we only finetune mLED with our training set of the same direction. We refer to this model as mLED-Base.
For baseline 2 (baseline 1 + historical machine translation), we utilize our historical translation dataset. We first finetune with the translation task and then with the target task, which we call mLED-Historical MT.
(3) Finally, for baseline 3, we finetune with datasets from both directions. We add prefixes to indicate the directions. We refer to this model as mLED-.
Intermediate task finetuning Apart from the final finetuning step with our target CLCTS task, we conduct intermediate task finetuning using external sources including MLS, CLS, and CTS. Table 6 depicts the detailed design.
@@
(2) Direction hEn-De |
4.1.3 CLCTS with ChatGPT
@@
via the OpenAI API to query the summary. We set the temperature to 0.7 to allow for some randomness. We adopt (1) a pipeline translate-then-summarize prompt as well as a (2) e2e prompt. We set prompt lines to the same language as the target output since ChatGPT works better in this setting according to our initial experiments.111111During our first round of data collection all the prompt lines were in English. We noticed that ChatGPT tends to output summaries in the wrong target language, especially when the summary is expected to be in German.
We truncate the texts that exceed ChatGPT input limits.121212According to our experiment with the GPT2 tokenizer, we find that on average our German texts have 1.73 tokens per word (which is higher than the suggested 1.3 tokens per word ratio from OpenAI) and English 1.07 tokens per word. Therefore, we limit the German text to 2,048 words and English to 3,000 words. Since the 4,096 tokens limits from GPT-3.5-turbo are shared between both prompts and completions (outputs), we need to leave some tokens for the output summary.
@@
Retrieve-then-summarize with ChatGPT To tackle input limitations of ChatGPT (e.g., for GPT-3.5-turbo a total of 4,096 tokens are shared between prompt and completion),131313https://help.openai.com/en/articles/4936856-what-are-tokens-and-how-to-count-them we build the retrieve-then-summarize design, where we use MemSum as a key sentence retriever and input the retrieved sentences to ChatGPT and request summarization in the target language. For this experiment, we limit the maximal extracted sentences per document to 100 according to the input limits of ChatGPT.
Type | Src-Tgt | Prompt |
e2e | hDe-En | Please summarize the following text in English : [Text]. |
e2e | hEn-De | Bitte fasse den folgenden Text auf Deutsch zusammen: [Text]. |
e2e (Title) | De-En | Please give me the summary of the story [Title] written by [Author]. |
e2e (Title) | En-De | Bitte gebe mir die Zusammenfassung der Geschichte [Titel] von [Autor]. |
pipeline | hDe-En | Please first translate the following text into English and summarize the translated text: [Text] |
pipeline | hEn-De | Bitte übersetze zuerst den folgenden Text auf Deutsch und fasse den übersetzten Text zusammen: [Text]. |
@@ 4.2 Evaluation
To identify the best model, we utilize multiple recently popular reference-based metrics. We also conduct a multi-phase human evaluation to evaluate the model and test the effectiveness of the resulting metrics. During the multi-phase evaluation, we give feedback to annotators after each phase for them to better align with the instructions. We list the details of all three evaluation strategies in Table 8.
Type | Aspects | Source | Detail |
Lexical overlap | ROUGE-variant Lin (2004) | ROUGE-1 matching the texts through computing the unigram overlap; ROUGE-L leveraging the longest common subsequence | |
Automatic | Semantic overlap | BERTScore Zhang et al. (2020b) | measuring semantic overlap through calculating the token similarity |
MoverScore Zhao et al. (2019) | measuring semantic overlap utilizing Word Mover’s distance | ||
Generation-based | BARTScore Yuan, Neubig, and Liu (2021) | metric based on text generation using BART | |
NLI-based | MENLI Chen and Eger (2023) | metric based on natural language inference | |
Discourse-based | DiscoScore Zhao, Strube, and Eger (2023) | metric using BERT to model discourse coherence | |
Human | consistency, coherence, fluency, and relevance | SummEval Fabbri et al. (2021) | 6 annotators, 130 annotated instances for direction hDe-En; 104 annotated instances for hEn-De. |
ChatGPT | prompting | - | prompts consisting of the same instruction for human annotators; 494 annotations for hDe-En & 407 for hEn-De. |
Automatic Evaluation
We consider six scores as automatic metrics to reflect five different aspects for measuring the quality of output summaries. The configurations used for each automatic metric are presented in Table 18 (appendix).
Human Evaluation We use six ‘expert’ annotators to evaluate all models. Three of the annotators are female and three are male. All annotators have high competence in English as they are affiliated with the university: they are Master students (3), postgraduates (1), PhD students (1), and NLP faculty members (1). Three of them are native German speakers. Thus, the generated English summaries are evaluated by all six annotators and the generated German summaries are annotated by three native speakers.
We utilize the evaluation strategy by Fabbri et al. (2021) where the annotators score the summary from four perspectives: coherence, consistency, fluency, and relevance. The outputs are rated with a score ranging from 1 (worst) to 5 (best) for each dimension and we allow 0.5 increments. The annotators are presented with the source document, one reference summary, and one output summary during annotation.141414The annotation details including instructions and the number of annotated instances per model are given in Section 9.5 (appendix). After phase 1, we give feedback to the annotators by providing a comparison of their scores with the average scores. The scores that differ substantially are highlighted and we encourage them to review the differences.
@@
Nonetheless, we make sure that each model is evaluated for at least 4 identical source documents. For hDe-En, we collect 130 summary level instances and 104 instances for hEn-De.
ChatGPT evaluation Since human annotation is costly, we follow related work Gao et al. (2023); Chiang and Lee (2023) and test ChatGPT as an evaluator for our CLCTS task. We differ in the following aspects: (1) our CLCTS task is harder to evaluate due to the long length of both reference and generated summary. (2) Our CLCTS task involves multilingual evaluation for both English and German. We give the same instructions to ChatGPT as to the human annotators. For reproducibility purposes, we set the temperature to 0 to reduce the randomness. We collect evaluations for all available generated summaries from documents that have been annotated by humans in the previous step, resulting in 494 ChatGPT annotations for hDe-En and 407 ChatGPT annotations for hEn-De.